# Machine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectra

### v1

@@ 9 Discussion
The validity of the results presented in this work may depend on the specific properties of the planet, its atmosphere, and the star we selected for our numerical experiments. In the following we will analyse the sensitivity of our results to those variable factors.
@@
Changing the radius of the star or the radius of the planet will only imply a change in the signal strength and noise and hence a different value the SNR. Pressure, which essentially determines the height scale of the atmosphere, will have an analogous effect. Since our experiments explore a large range of SNR, any modification of those parameters will fall in one of the synthetic noisy spectra we use for training and testing. In other words, a variation in SNR is equivalent to a variation in either the ratio of the radius of the planet to the radius of the star or the atmospheric pressure. Of course, in real atmospheres, all these factors may affect the equilibrium composition of the atmosphere. However, the model atmospheres we use in this work, that in some sense is a demonstrative effort, are simple enough in such a way that chemistry will not be related to these factors.
Changing the spectral type of the star or the temperature of the planetary atmosphere would not only change the SNR but also the transmission spectra in non-trivial way. This changes may actually impact the performance of the algorithms.
To test this, we performed additional numerical experiments using, first, different stellar effective temperatures in the range of M and K-dwarf stars. We observe that changing the stellar spectrum does not significantly affect the classification results.
In the second place, the effect of atmospheric temperature was implicit to our experiments. Actually, and as described, for instance, in subsection 6.1, our training and test sets already include transmission spectra with different atmospheric temperatures (in the range of 200 K to 400 K).
@@
We observe that at lower atmospheric temperatures the performance of the algorithms improves. For instance, the minimum mixing ratio for detection, at the lowest atmospheric temperatures (200 K), may be reduced by one order of magnitude at very low values of the SNR, for several of the interesting molecules. The reason for this temperature dependence is the change in the molecular cross-section that affects the capability of the algorithm to recognize the presence of the molecules.
In summary, our models are robust against variations in stellar temperature and stellar radius, at least in the range of spectral types considered in this work (K and M dwarfs). However, our experiments reveal that planetary atmospheric temperature remains a factor that may influence the accuracy of biosignature classification.
@@
Since the performance of the SC algorithms depends on the structure of the noise, one question that can be raised is why we do not train our algorithms using the specific noise of NIRSpec. Here, we should recall that our work was aimed to assess the feasibility of using machine learning to classify planets with biosignatures with low SNR, but not with a specific noise structure. If we have chosen a certain type of noise, the validity of our results could become noise-dependent and the generality of our conclusions will be lost. Still, the issue of studying the impact that the noise structure has on the performance of classification algorithms should be further investigated.
@@ 10 Summary and conclusions
In this work, we explored the use of supervised machine learning algorithms to classify low SNR exoplanet transmission spectra, based on their potential for containing biosignatures. We employed Random Forest (RF) for designing and training three types of models: a binary classificator (BC) which labels a planet according to whether it is interesting or not for follow-up observations; a multilabel classificator (MC) which labels an observed spectrum according to whether it may have certain molecular species; and specialised classificators (SC) that work similar to MC by assembling multiple binary RF classificators focused on particular molecules.
For training and testing the algorithms, we generated of the order of $10^{6}$ synthetic noisy spectra, corresponding to a planet with the same physical characteristics as TRAPPIST-1 e, but having a wide diversity of compositions and at three atmospheric temperatures. For that purpose, we designed a package, MultiREx [^3] http://pypi.org/project/multirex, that works as a wrapping for TauREx (Al-Refaie et al., 2021) but has the capabilities of generating, representing and manipulating large sets of synthetic noisy spectra, especially intended for machine learning purposes.
@@
We tested our models by measuring different binary and multilabel standard metrics (Recall, TNR, F1 Score, etc.) that we interpret, in the context of biosignature searching, as time-saving metrics (minimising the time spent on following-up not interesting planets), wasting metrics (not excluding inhabited planets as non-interesting candidates), discovery metrics (maximising the opportunity to achieve an actual discovery), among others.
@@
The BC algorithm, although lacking the accuracy for detailed molecular classification, achieves the highest values of the discovery metric. With observed spectra having values of the SNR as low as 1 or 2, which can be achieved, according to our estimations, by observing 1-5 transits with the JWST NIRSpec PRISM, the BC model labels as interesting most of the inhabited planets in a survey sample, even if they have a huge range of biosignatures mixing-ratios. Moreover, the algorithm performs incredibly well (TNR$>80\%$) at tagging as not interesting, planets without biosignatures, avoiding wasting time on follow-up observations.
@@
Our results show that if we perform a survey of potentially habitable exoplanets and measure the transmission spectra using only 1 to 5 transits, we can identify most of the inhabited exoplanets in the survey for performing follow-up, time-intensive observations, only if their atmospheres have surface mixing ratios of CH4 similar to those of present and prebiotic Earth ($\mathcal{O}(10^{-6})$, Kaltenegger et al. 2007, 2020). This number of transits is significantly lower than those required for the detection of the same molecule using standard and non-standard retrieval procedures.
It is important to stress that the methodology proposed here is not intended for detecting biosignatures or for replacing them. In this sense, all the efforts intended at performing full retrieval are still required. Identifying a planet as interesting will only make the allocation of observing time of valuable resources such as JWST more efficient, which is an important goal in modern astronomy.
Classifying planets based on the presence of ozone will be harder than in the case of methane. With mixing ratios of the order of $10^{-7}$, which are slightly lower than those of present Earth, efficient classification with our algorithms is possible for values of the SNR of 5 (SC) and 6 (MC), which corresponding to 30-50 transits. Although this number of transits seems to be large, it is important to recall that standard retrieval procedures for ozone at mixing ratios as high as $10^{-6}$ require more than 100 transits Lin et al. (2021). On the other hand, for planets orbiting cool M-dwarf stars, like TRAPPIST-1 e, higher mixing ratios of O3 with lower levels of O2 could be common if the star’s UV emission promotes its production (Grenfell et al., 2014).
@@
For future research, it would be valuable to explore the integration of other machine learning methods and increase the diversity of the synthetic spectra used for training. Additionally, including other potential biosignatures and atmospheric conditions could improve the capacity to generalise the models. These advances will contribute to more accurate and efficient biosignature detection, supporting ongoing efforts in the search for life beyond our Solar System.
11 Data availability and reproducibility

### v4

@@ 10 Discussion
The validity of the results presented in this work may depend on the specific properties of the planet, its atmosphere, and the star selected for our numerical experiments. In the following, we analyse the sensitivity of our results to these variable factors.
@@
If we disregard the effect of atmospheric refraction (see below), changing the radius of the star or the planet simply alters the signal strength and noise, and thus the SNR value. Pressure, which essentially determines the atmospheric scale height, has a similar effect. Since our experiments span a broad range of SNR values, any change in these parameters will lie within one of the synthetic noisy spectra used for training and testing. In other words, a variation in SNR is equivalent to a variation in either the planet-to-star radius ratio or the atmospheric pressure. Although these factors influence equilibrium compositions in realistic atmospheric models, the simplified model atmospheres employed in this study do not explicitly link atmospheric chemistry to these parameters. We emphasise that our aim is to demonstrate the feasibility of this approach, rather than to design robust algorithms for an actual observing campaign.
Other studies (e.g. Bétrémieux & Kaltenegger 2014) have shown that atmospheric refraction limits how deeply an exoplanet’s atmosphere can be probed, and that this limitation depends on the host star’s radius. For M-dwarf stars—which appear smaller from the perspective of a habitable-zone planet (see Figure 2 in Bétrémieux & Kaltenegger 2014)—the refractive effect is minimal, and stellar light can reach very deep atmospheric layers, sometimes down to the surface. Conversely, for K-dwarf stars, the larger angular size results in a higher critical altitude. For instance, in the case of a K0 star, stellar light may only penetrate to about 8–14 km, which, for present-day Earth, lies above the troposphere where methane and water are most abundant.
Although our model does not explicitly include these refractive effects, the results of such studies allow us to affirm generally that the host star’s spectral type influences the accessible atmospheric depth—minimal for M-dwarfs but more significant for K-dwarfs, particularly if clouds are present.
Changing the star’s effective temperature (and thus its spectral energy distribution) or the planetary atmospheric temperature would not only affect the SNR but also modify the transmission spectra in a non-trivial manner, potentially impacting algorithm performance. To test the effect of varying stellar spectra, we conducted additional numerical experiments with different stellar effective temperatures in the M- and K-dwarf ranges. We find that altering the stellar spectrum does not significantly affect classification results.
@@
On the other hand, the effect of atmospheric temperature was already implicit in our experiments. As described, for instance, in subsection 7.1, our training and test sets include transmission spectra at various atmospheric temperatures (from 200 K to 400 K). We observe that at lower temperatures, algorithm performance improves. For example, at 200 K, the minimum mixing ratio required for detection can be reduced by an order of magnitude at very low SNR values for several molecules. This temperature dependence arises because molecular cross-sections—especially that of CO2—change with temperature, affecting the algorithm’s ability to recognise molecular signatures.
In summary, our models are robust against variations in stellar temperature and radius, at least for M-dwarfs. The case of K-dwarfs should be examined independently to account for atmospheric refraction. We anticipate that this effect would merely shift the mixing ratio thresholds obtained here without altering the main conclusions of this work. Our experiments indicate that planetary atmospheric temperature can affect the accuracy of biosignature/bioindicator classification.
@@
Since the performance of the SC algorithms depends on the noise structure, one might ask why we did not train our algorithms using the specific noise profile of NIRSpec. We reiterate that our objective was to assess the feasibility of using machine learning to classify planets with biosignatures/bioindicators at low SNR, rather than optimising for a particular noise structure. Focusing on a specific type of noise could render our results noise-dependent and undermine the generality of our conclusions. Nevertheless, further investigation of how the noise structure affects classification algorithm performance is warranted.
@@ 11 Summary and conclusions
In this work, we explored the use of supervised machine learning algorithms to classify low-SNR exoplanet transmission spectra based on their potential for containing biosignatures/bioindicators. We employed Random Forest (RF) to design and train three types of models: a binary classifier (BC) that labels a planet as interesting or not for follow-up observations; a multilabel classifier (MC) that assigns labels to an observed spectrum according to the presence of certain molecular species; and specialised classifiers (SC) that operate similarly to MC by assembling multiple binary RF classifiers focused on particular molecules.
For training and testing, we generated on the order of $10^{7}$ synthetic noisy spectra, including stellar contamination, corresponding to a planet with physical characteristics similar to TRAPPIST-1 e, but covering a wide diversity of compositions and three atmospheric temperatures.
Generating and handling such a large set of spectra is challenging. For this purpose, we developed a package, MultiREx [^4] http://pypi.org/project/multirex, which serves as a wrapper for TauREx (Al-Refaie et al., 2021) and is specifically designed to generate, represent, and manipulate large sets of synthetic noisy spectra for machine learning applications.
@@
We tested our models by measuring various binary and multilabel standard metrics (Recall, TNR, F1 Score, etc.), which we interpret in the context of biosignature/bioindicator searches as time-saving metrics (minimising the time spent on following up uninteresting planets), wasting metrics (ensuring inhabited planets are not erroneously excluded), discovery metrics (maximising the opportunity for actual discovery), among others.
@@
Although the BC algorithm lacks the precision for detailed molecular classification, it achieves the highest values for the discovery metric (F1 Score). With observed spectra having SNR values as low as 1 or 2—which, according to our estimations, can be achieved by observing 1–5 transits with the JWST NIRSpec PRISM—the BC model labels most of the inhabited planets in a survey sample as interesting, even when they exhibit a wide range of biosignature/bioindicator mixing ratios. Moreover, the algorithm performs exceptionally well (TNR > 80%) at tagging planets without biosignatures/bioindicators as not interesting, thereby avoiding wasted follow-up observations.
@@
Our results indicate that by surveying potentially habitable exoplanets with transmission spectra obtained from only 4 to 10 transits, we can identify most inhabited exoplanets for follow-up, time-intensive observations—provided that their atmospheres have integrated mixing ratios of CH4 and O3 similar to those of present and prebiotic Earth ($\mathcal{O}(10^{-6})$ and $\mathcal{O}(10^{-5})$, respectively; Lustig-Yaeger et al. 2023; Lin et al. 2021). This number of transits is significantly lower than that required by standard and non-standard retrieval procedures for detecting these molecules.
It must be emphasised that the methodology proposed here is not intended to replace more detailed retrieval methods or to directly detect biosignatures/bioindicators. Rather, it aims to enhance the efficiency of observing time allocation for valuable resources such as JWST by identifying interesting targets for follow-up observations—an important goal in modern astronomy.
@@
For future research, it would be valuable to explore the integration of additional machine learning methods and to increase the diversity of the synthetic spectra used for training. Furthermore, incorporating other potential biosignatures or bioindicators, along with a broader range of atmospheric conditions, could enhance the models’ generalisability. These advances will contribute to more accurate and efficient detection of biosignatures/bioindicators, thereby supporting ongoing efforts in the search for life beyond our Solar System.
Acknowledgments
We sincerely acknowledge the referee for their insightful comments and valuable suggestions, especially for proposing additional aspects to evaluate and train our models, which significantly enhance their robustness.
Data availability and reproducibility